OpenStudy (anonymous):

please help solve this a and b are the legs of a rectangle triangle if we know the hypotenuse is 4 times leg a then the the area of the triangle in terms of b would be??

OpenStudy (jdoe0001):

a rectangle triangle ?

OpenStudy (jdoe0001):



OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

(4a)^2=a^2+b^2 16a^2=a^2+b^2 16a^2-a^2=b^2 15a^2=b^2 √15a^2=b b=a√15 > so i got b